# QuiZSF: An efficient data-model interaction framework for zero-shot time-series forecasting

### v1

@@
Abstract
Time series forecasting has become increasingly important to empower diverse applications with streaming data. Zero-shot time-series forecasting (ZSF), particularly valuable in data-scarce scenarios, such as domain transfer or forecasting under extreme conditions, is difficult for traditional models to deal with. While time series pre-trained models (TSPMs) have demonstrated strong performance in ZSF, they often lack mechanisms to dynamically incorporate external knowledge. Fortunately, emerging retrieval-augmented generation (RAG) offers a promising path for injecting such knowledge on demand, yet they are rarely integrated with TSPMs. To leverage the strengths of both worlds, we introduce RAG into TSPMs to enhance zero-shot time series forecasting. In this paper, we propose QuiZSF (Quick Zero-Shot Time Series Forecaster), a lightweight and modular framework that couples efficient retrieval with representation learning and model adaptation for ZSF. Specifically, we construct a hierarchical tree-structured ChronoRAG Base (CRB) for scalable time-series storage and domain-aware retrieval, introduce a Multi-grained Series Interaction Learner (MSIL) to extract fine- and coarse-grained relational features, and develop a dual-branch Model Cooperation Coherer (MCC) that aligns retrieved knowledge with two kinds of TSPMs: Non-LLM based and LLM based. Compared with contemporary baselines, QuiZSF, with Non-LLM based and LLM based TSPMs as base model, respectively, ranks Top1 in 75% and 87.5% of prediction settings, while maintaining high efficiency in memory and inference time.
1 Introduction
Time series forecasting (TSF) is crucial for data-driven applications in dynamic systems, ranging from weather prediction, electricity consumption estimation to economic dynamic forecasting. Despite the excellent performance of many TSF studies, they face challenges in data-scarce situations wu2022timesnet ; gruver2024large ; jin2023time , such as extreme weather prediction and cross-domain adaptation, where limited historical data hinder accurate forecasting. Since it is impossible to summarize all series patterns within a system liutimer , zero-shot time forecasting (ZSF), which exploits available patterns to infer the trends and future evolution of new time series, has become a prevailing and essential research topic ekambaram2024tiny .
Recently, there has been growing interest in developing general times series pre-trained models (TSPMs) for ZSF. These models aim to transfer learned knowledge to previously unseen target time series (TS) datasets, similar to the success achieved in natural language processing (NLP) and computer vision (CV). TSPMs can be mainly classified into two categories: Non-LLM based TSPMs and LLM based TSPMs (ekambaram2024tiny ). The former pre-trains models on multiple TS datasets, represented by models like Moment (goswami2024moment ), TTM (ekambaram2024tiny ), TimesFM (das2024decoder ), Moirai (woo2024unified ), and Lag-LLaMA (rasul2023lag ); the latter uses pre-trained large language models (LLMs) for TSF, treating the prediction task as a form of cross-domain transfer learning, represented by models such as LLMTime (gruver2024large ), Time-LLM (jin2023time ), and GPT4TS (zhou2023one ). Both types of TSPMs have demonstrated good performance in ZSF.
However, using TSPMs for ZSF faces two main issues: (1) Real-world data is highly dynamic, but TSPMs cannot update their knowledge in real-time. Fine-tuning on new data is computationally expensive and disrupts the plug-and-play goal of ZSF tan2024language . The issue is how to enable models to learn dynamically updated knowledge without incurring excessive costs while maintaining a plug-and-play approach. (2) Time series often exhibit structural similarity tire2024retrieval , which is especially valuable in sparse-data ZSF settings. The issue is how to leverage inter- and intra-domain data with patterns similar to the target sequence to enhance prediction accuracy.
To address these issues, we propose integrating TSPMs with retrieval-augmented generation (RAG). For issue (1), RAG provides efficient access to updated external knowledge, avoiding costly fine-tuning and preserving the plug-and-play nature of ZSF. For issue (2), RAG retrieves relevant sequences from inter- and intra-domain datasets, offering contextual information such as periodic trends, spikes, or transitions. This enhances model generalization and prediction accuracy in sparse-data scenarios. As shown in Figure 1, RAG identifies similar TS in other datasets to improve ZSF for ETTm1.
@@ 1 Introduction
Figure 1: Motivation. Time series across domains often exhibit similar temporal patterns, which can be retrieved and reused as auxiliary knowledge.
@@
Nevertheless, integrating external retrieval with TSFMs introduces three fundamental challenges: (i) Efficient and Domain-Sensitive Storage and Retrieval: How to structure and index millions of TS for fast, relevant, and cross-domain retrieval remains a foundational bottleneck in large-scale systems. (ii) Multi-Level Feature Extraction: Retrieved sequences vary in scale, domain, and noise level. Designing a lightweight yet expressive architecture to extract multi-level interaction patterns between target and retrieved series is essential for generalization. (iii) Modality-Aligned Representation Integration: TSPMs are categorized into Non-LLM based and LLM based types. The former requires seamless fusion of numerical features, while the latter demands coherent transformation of retrieved data into language prompts. A unified adapter must bridge these modalities while supporting gradient-based optimization.
@@
To address these challenges, we propose the Quick Zero-Shot Time Series Forecaster (QuiZSF), which, to the best of our knowledge, is the first efficient data–model interaction framework for ZSF. QuiZSF integrates scalable storage and retrieval, informative sequence representation learning, and modality-aware model adaptation. Our main contributions are summarized as follows:
@@
  • We construct the ChronoRAG Base (CRB), a hierarchical tree-structured temporal database, and propose a Hybrid and Hierarchical TS Retrieval (HHTR) strategy that enables fast, scalable, and domain-sensitive retrieval across large-scale time-series databases.
@@
  • We propose the Multi-grained Series Interaction Learner (MSIL), a lightweight yet expressive module that extracts both fine-grained dependencies and coarse-grained patterns from retrieved sequences to enhance target sequence understanding.
@@
  • We present the Model Cooperation Coherer (MCC), a dual-branch adapter that integrates retrieved features into both Non-LLM and LLM-based TSPMs. MCC constructs residual-semantic fusion inputs for numerical models and structured prompts for language models, while supporting end-to-end optimization across modalities.
@@
  • We conduct experiments using Non-LLM and LLM based TSPMs as base models, separately. The former achieves first place in 75% of ZSF settings, while the latter ranks first in 87.5% of the settings, all while maintaining high efficiency in memory usage and inference speed.
2 Related Work
2.1 TSPMs for Zero-shot Time Series Forecasting
@@
Recent advancements in TSPMs for ZSF have garnered significant attention. These models can be broadly categorized into two types. The first is pre-trained models designed specifically for TSF. These models learn generalizable temporal representations from large-scale TS datasets.Representative methods include Moment goswami2024moment , TTM ekambaram2024tiny , TimesFM das2024decoder , Moirai woo2024unified , and Lag-LLaMA rasul2023lag . For example, TTM adopts a lightweight mixer-style architecture and demonstrates strong zero-shot performance across multiple domains. The second type leverages pre-trained large language models (LLMs) by framing time-series forecasting as a form of cross-domain transfer learning. These models, including LLMTime gruver2024large , Time-LLM jin2023time , and GPT4TS zhou2023one , transform numerical sequences into natural language prompts, enabling ZSF.
@@
Despite promising results, these models rely heavily on large-scale pre-training or fine-tuning across diverse datasets. They often lack scalability and the ability to incorporate real-time, open-world knowledge. As illustrated in Figure 1, time series from similar domains frequently share common temporal patterns, which can serve as valuable references in low-resource settings. With the advancement of retrieval techniques, integrating external sequence-level knowledge into pre-trained models enables dynamic forecasting, mitigates hallucinations, and constrains predictions within realistic boundaries. Enhancing TSPMs with retrieval capabilities has therefore emerged as a promising direction for improving zero-shot forecasting.
2.2 Retrieval Augmented Generation
@@
Retrieval-Augmented Generation (RAG) combines models with information retrieval to enhance model performance by leveraging external knowledge. Early work, such as REALM guu2020retrieval , introduces retrieval-augmented models by incorporating unsupervised retrieval modules, allowing models to access relevant passages during generation. RAG lewis2020retrieval further advances this by combining retrieval and generation in an end-to-end framework, improving tasks like open-domain question answering. Subsequent approaches, including T5+RAG lewis2020retrieval and Fusion-in-Decoder (FiD) izacard2020leveraging , enhance the fusion of retrieved information, achieving better coherence and relevance. Models like RETRO borgeaud2022improving have pushed the boundaries by incorporating retrieval at both pre- and post-generation stages, further improving output quality. Applying RAG to time series forecasting offers promising benefits. Recent studies jing2022retrieval ; yeh2023efficient ; tire2024retrieval have introduced retrieval mechanisms into forecasting. However, they often overlook the cost and scalability of external databases in large-scale settings, leading to inefficiencies. While TimeRAF zhang2024timeraf explores zero-shot forecasting with retrieval, it lacks a clear mechanism for modeling interactions between pre-trained models and retrieved sequences. These pioneering studies highlight the potential value of combining pre-trained models with retrieval processes.
3 Preliminaries and Problem Definition
@@
3.1 Retrieval-Augmented Non-LLM based TSPMs
@@
3.2 Retrieval-Augmented LLM based TSPMs
@@
$\widehat{\mathbf{X}}^{T},\mathcal{P}=\mathcal{F}_{\text{L}}^{*}(\mathbf{X}^{T},\mathbb{X}^{R}),\quad\widehat{\mathbf{Y}}^{T}=\mathcal{M}_{2}^{*}(\widehat{\mathbf{X}}^{T}\mid\mathcal{P}),$ (1)
@@
where $\mathcal{M}_{2}$ is the modified learning framework with LLM based TSPM and $\mathcal{P}$ is the generated prompt fusing the retrieved knowledge.
Figure 2: Overview of QuiZSF.
4 Methodology
4.1 Framework overview
@@
Our zero-shot learning framework consists of three major components, as illustrated in Figure 2. 1) a ChronoRAG Base with efficient hierarchical retrieval, 2) a Multi-grained Series Interaction Learner captures series-level interactions between the target and retrieved series, offering valuable information for zero-shot forecasting, 3) a Model Cooperation Coherer enables modality-aligned integration between learned representations and both Non-LLM based and LLM based TSPMs.
4.2 ChronoRAG Base
@@
ChronoRAG Base (CRB) is constructed from 27 time series datasets across seven domains: Energy, Health, IoT, Nature, Web, Transport, and Environment. It integrates five public data sources: Monash godahewa2021monash , UCR dau2019ucr , TSER tan2021time , TDBrain van2022two , and UTSD liutimer . These datasets span diverse sampling granularities, ranging from daily to minute-level resolutions. We provide three CRB versions: Small, Medium, and Large, containing 34M, 48M, and 143M time points, respectively. All versions cover the same domains, and smaller versions are strict subsets of the larger ones. Detailed statistics of CRB-Large are shown in Table 5. Two core components support CRB as follows.
@@
Data protocol for ChronoRAG. To ensure consistency and scalability, we design a unified data protocol including sliding-window segmentation, linear interpolation for missing values, channel-independent processing for dimensional alignment, and standardized metadata. All sequences are stored in the ARROW format peltenburg2019fletcher for efficient access. Further Details are provided in Appendix A.1.
Hierarchical Series Tree. We implement a hierarchical tree structure to support fast indexing and retrieval. The base is partitioned by domain and clustered using $k$-means to enable efficient approximate search and incremental updates. Further Details are provided in Appendix 6.
4.3 Hybrid and Hierarchical Time-series Retrieval

### v2

# QuiZSF: A Retrieval-Augmented Framework for Zero-Shot Time Series Forecasting

@@
Abstract.
Accurate forecasting of sequential data streams is a cornerstone of modern Web services, supporting applications such as traffic management, user behavior modeling, and online anomaly prevention. However, in many Web environments, new domains emerge rapidly and labeled history data is scarce, which makes zero-shot forecasting particularly challenging. Existing time-series pre-trained models (TSPMs) show promise but they lack the ability to dynamically incorporate external knowledge, while conventional retrieval-augmented generation (RAG) methods are rarely extended beyond text. In this work, we present QuiZSF, a retrieval-augmented forecasting framework that integrates search and forecasting for time series data. The framework performs search by retrieving structurally similar sequences from a large-scale time-series database, and it performs forecasting by integrating the retrieved knowledge into the target sequence. Specifically, QuiZSF introduces a ChronoRAG Base, a hierarchical tree-structured database that enables scalable and domain-aware retrieval, a Multi-grained Series Interaction Learner that captures fine- and coarse-grained dependencies between target and retrieved sequences, and a Model Cooperation Coherer that adapts retrieved knowledge to TSPMs. This design teaches models to actively perform search, align auxiliary information across modalities, and leverage it for more accurate forecasting. Extensive experiments on five public benchmarks demonstrate that QuiZSF consistently outperforms strong baselines, ranking first in up to 87.5% of zero-shot forecasting settings while maintaining high efficiency.
Time Series Forecasting, Retrieval-Augmented Generation, Information Retrieval
$c$s: Computing methodologies → Artificial intelligence.
1. Introduction
Modern Web services rely heavily on sequential data streams, ranging from traffic monitoring and user behavior modeling to anomaly prevention in large-scale online platforms. Accurate forecasting of such time series is therefore a cornerstone for building reliable and intelligent Web applications. However, Web environments are highly dynamic and data-scarce: new domains and conditions frequently emerge, while historical labels are limited or unavailable. These challenges make zero-shot time-series forecasting (ZSF), which aims to infer future trends of unseen sequences without domain-specific supervision, a crucial yet difficult problem (Wu et al., 2022; Gruver et al., 2024; Jin et al., 2023; Ekambaram et al., 2024).
Recent progress in time series pre-trained models (TSPMs) has shown promising results for ZSF by transferring knowledge across datasets, analogous to pre-training in NLP and CV. TSPMs can be categorized into two groups (Ekambaram et al., 2024): (i) numerical Non-LLM based TSPMs such as Moment (Goswami et al., 2024), TTM (Ekambaram et al., 2024), TimesFM (Das et al., 2024), and Moirai (Woo et al., 2024), and (ii) textual LLM based models that reconstruct language models for time series forecasting, such as LLMTime (Gruver et al., 2024), Time-LLM (Jin et al., 2023), and GPT4TS (Zhou et al., 2023). Although both categories achieve strong forecasting accuracy, they face two critical limitations. First, real-world sequences are continually evolving, yet TSPMs cannot efficiently incorporate new knowledge without costly fine-tuning (Tan et al., 2024). Second, time series often exhibit structural similarity across domains (Tire et al., 2024), but TSPMs lack mechanisms to retrieve and reuse such auxiliary patterns to improve generalization.
To address these limitations, we propose integrating TSPMs with retrieval-augmented generation (RAG), a paradigm that has shown great success in natural language processing but remains underexplored for time series forecasting. In this approach, a search module first retrieves relevant sequences from a large-scale temporal database, and a forecasting module then integrates the retrieved knowledge into the target sequence prediction process. As illustrated in Figure 1, retrieved auxiliary sequences can provide valuable contextual signals such as periodicity, seasonal shifts, or abrupt transitions, which enhance forecasting quality and reduce hallucinations in zero-shot scenarios.
@@ 1. Introduction
Figure 1. Motivation. Time series across domains often exhibit similar temporal patterns, which can be retrieved and reused as auxiliary knowledge.
@@
Nevertheless, combining retrieval with forecasting introduces three core challenges: (i) Efficient and domain-sensitive storage and retrieval: indexing millions of sequences for fast and relevant cross-domain search remains difficult at Web scale; (ii) Multi-level feature extraction: retrieved sequences vary in scale, domain, and noise, requiring expressive yet lightweight interaction modeling; (iii) Modality-aligned representation integration: numerical Non-LLM based TSPMs demand feature-level fusion, while textual LLM-based models require structured prompts that translate retrieved sequences into language-compatible inputs.
@@
To address these challenges, we propose the Quick Zero-shot time-series Search and Forecasting framework (QuiZSF). QuiZSF is a retrieval-augmented forecasting framework that integrates scalable storage, cross-series interaction learning, and modality-aware adaptation. Our main contributions are summarized as follows:
@@
  • We build the ChronoRAG Base (CRB), a hierarchical tree-structured temporal database, together with a Hybrid and Hierarchical TS Retrieval (HHTR) strategy for fast and domain-sensitive search.
@@
  • We design the Multi-grained Series Interaction Learner (MSIL), a module that captures fine-grained dependencies and coarse-grained trends from retrieved sequences to improve target sequence understanding.
@@
  • We develop the Model Cooperation Coherer (MCC), a dual-branch adapter that integrates retrieved information into both Non-LLM and LLM-based TSPMs.
@@
  • We conduct extensive experiments on public benchmarks, where QuiZSF outperforms state-of-the-art baselines. Using Non-LLM TSPMs, QuiZSF achieves top performance in 75% of ZSF settings; with LLM-based TSPMs, it ranks first in 87.5% of settings, while maintaining high efficiency in memory usage and inference speed.
Figure 2. Overview of QuiZSF.
2. Related Work
2.1. TSPMs for Zero-shot Forecasting
@@
Recent advancements in TSPMs for ZSF have garnered significant attention. These models can be broadly categorized into two types. The first is pre-trained models designed specifically for TSF. These models learn generalizable temporal representations from large-scale TS datasets.Representative methods include Moment (Goswami et al., 2024), TTM (Ekambaram et al., 2024), TimesFM (Das et al., 2024), Moirai (Woo et al., 2024), and Lag-LLaMA (Rasul et al., 2023). The second type leverages pre-trained large language models (LLMs) by framing time-series forecasting as a form of cross-domain transfer learning. These models, including LLMTime (Gruver et al., 2024), Time-LLM (Jin et al., 2023), and GPT4TS (Zhou et al., 2023), transform numerical sequences into natural language prompts, enabling ZSF.
@@
Despite promising results, these models rely heavily on large-scale pre-training or fine-tuning across diverse datasets. They often lack scalability and the ability to incorporate real-time, open-world knowledge. As illustrated in Figure 1, time series from similar domains frequently share common temporal patterns, which can serve as valuable references in low-resource settings. With the advancement of retrieval techniques, integrating external sequence-level knowledge into pre-trained models enables dynamic forecasting, mitigates hallucinations, and constrains forecasting within realistic boundaries. Enhancing TSPMs with retrieval capabilities has therefore emerged as a promising direction for improving zero-shot time-series forecasting.
2.2. Retrieval Augmented Generation
@@
Retrieval-Augmented Generation (RAG) combines models with information retrieval to enhance model performance by leveraging external knowledge. Early work, such as REALM (Guu et al., 2020), introduces retrieval-augmented models by incorporating unsupervised retrieval modules, allowing models to access relevant passages during generation. RAG (Lewis et al., 2020) further advances this by combining retrieval and generation in an end-to-end framework, improving tasks like open-domain question answering. Subsequent approaches, including T5+RAG (Lewis et al., 2020) and Fusion-in-Decoder (FiD) (Izacard and Grave, 2020), enhance the fusion of retrieved information, achieving better coherence and relevance. Applying RAG to time series forecasting offers promising benefits. Recent studies (Jing et al., 2022; Yeh et al., 2023; Tire et al., 2024) have introduced retrieval mechanisms into forecasting. However, they often overlook the cost and scalability of external databases in large-scale settings, leading to inefficiencies. While TimeRAF (Zhang et al., 2024) explores zero-shot forecasting with retrieval, it lacks a clear mechanism for modeling interactions between pre-trained models and retrieved sequences. These pioneering studies highlight the potential value of combining pre-trained models with retrieval processes.
3. Preliminaries and Problem Definition
@@
3.1. Retrieval-Augmented Non-LLM based TSPMs
@@
3.2. Retrieval-Augmented LLM based TSPMs
@@
(1) $\widehat{\mathbf{X}}^{T},\mathcal{P}=\mathcal{F}_{\text{L}}^{*}(\mathbf{X}^{T},\mathbb{X}^{R}),\quad\widehat{\mathbf{Y}}^{T}=\mathcal{M}_{2}^{*}(\widehat{\mathbf{X}}^{T}\mid\mathcal{P}),$
@@
where $\mathcal{M}_{2}$ is the modified learning framework and $\mathcal{P}$ is the generated prompt fusing the retrieved knowledge.
4. Methodology
4.1. Framework overview
@@
Our zero-shot learning framework consists of three major components, as illustrated in Figure 2. (i) a ChronoRAG Base with efficient hierarchical retrieval, (ii) a Multi-grained Series Interaction Learner captures series-level interactions between the target and retrieved series, offering valuable information for zero-shot forecasting, (iii) a Model Cooperation Coherer enables modality-aligned integration between learned representations and both Non-LLM based and LLM based TSPMs.
4.2. ChronoRAG Base
@@
ChronoRAG Base (CRB) comprises 27 time series datasets categorized into seven domains: Web, Energy, Health, IoT, Nature, Transport, and Environment. It primarily integrates five open data sources: UTSD (Liu et al., ), TSER Archive (Tan et al., 2021), Monash (Godahewa et al., 2021), TDBrain (Van Dijk et al., 2022), and UCR Time Series Archive (Dau et al., 2019). These datasets exhibit diverse sampling frequencies, ranging from macro-level intervals (e.g., daily) to finer granularities (e.g., hourly or minute-level). Notably, some datasets demonstrate exceptionally high sampling rates, such as the TDBrain dataset, which operates at a millisecond-level frequency. Based on these datasets, we constructed three versions of the database: CRB-Small, CRB-Medium, and CRB-Large, containing 34M, 48M, and 143M time points, respectively. Detailed information about CRB-Large can be found in Table 1. Each database covers all seven domains, and the smaller versions are subsets of the larger ones. There are three important database designs, include a well-designed data protocol for ChronRAG for processing high-quality datasets, a Hierarchical Series Tree for efficient series-level storage as well as a Hybrid and Hierarchical Time-series Retrieval for efficient retrieval.
Table 1. CRB-Large Detailed Descriptions: Domain indicates the field to which the dataset belongs. Datasets refer to the specific datasets included. Time Series represents the number of time series contained in the dataset after processing. Frequency denotes the sampling interval of time points, where “-” indicates either the absence of timestamps or irregular intervals. Time Points represents the total number of time points in the dataset. Source specifies the original paper or resource from which the dataset is obtained.
Domain Datasets Time Series Frequency Time Points Source
Web kaggle_web_traffic_dataset_without_missing_values 141444 Daily 72419328 Monash (Godahewa et al., 2021)
Energy wind_4_seconds_dataset 1 4 Sec 512 Monash (Godahewa et al., 2021)
australian_electricity_demand_dataset 5 30 Min 2560 Monash (Godahewa et al., 2021)
london_smart_meters_dataset_without_missing_values 5556 Hourly 2844672 Monash (Godahewa et al., 2021)
Health SelfRegulationSCP1 3366 0.004 Sec 1723392 UCR Time Series Archive (Dau et al., 2019)
MotorImagery 24192 0.001 Sec 12386304 UCR Time Series Archive (Dau et al., 2019)
PigCVP 312 - 159744 UCR Time Series Archive (Dau et al., 2019)
PigArtPressure 312 - 159744 UCR Time Series Archive (Dau et al., 2019)
SelfRegulationSCP2 2660 0.004 Sec 1361920 UCR Time Series Archive (Dau et al., 2019)
AtrialFibrillation 60 0.008 Sec 30720 UCR Time Series Archive (Dau et al., 2019)
IEEEPPG 15480 0.008 Sec 7925760 TSER archive (Tan et al., 2021)
BIDMC32HR 12278 - 6286336 TSER archive (Tan et al., 2021)
TDBrain 28644 0.002 Sec 14665728 TDBrain (Van Dijk et al., 2022)
IoT baian 918 0.02 Sec 470016 UTSD  (Liu et al., )
Nature StarLightCurves 9236 - 4728832 UCR Time Series Archive (Dau et al., 2019)
Phoneme 2110 - 1080320 UCR Time Series Archive (Dau et al., 2019)
EigenWorms 1554 - 795648 UCR Time Series Archive (Dau et al., 2019)
Worms 258 0.033 Sec 132096 UCR Time Series Archive (Dau et al., 2019)
us_births_dataset 1 Daily 512 Monash (Godahewa et al., 2021)
kdd_cup_2018_dataset_without_missing_values 270 Hourly 138240 Monash (Godahewa et al., 2021)
temperature_rain_dataset_without_missing_values 32072 Daily 16420864 Monash (Godahewa et al., 2021)
Sunspot_dataset_without_missing_values 1 Daily 512 Monash (Godahewa et al., 2021)
saugeenday_dataset 1 Daily 512 Monash (Godahewa et al., 2021)
Transport pedestrian_counts_dataset 66 Hourly 33792 Monash (Godahewa et al., 2021)
Environment AustraliaRainfall 3 Hourly 1536 TSER archive (Tan et al., 2021)
BenzeneConcentration 8 Hourly 4096 TSER archive (Tan et al., 2021)
BeijingPM25Quality 9 Hourly 4608 TSER archive (Tan et al., 2021)
4.2.1. Data protocol for ChronoRAG.
@@
To ensure consistency and scalability, we design a unified data protocol including sliding-window segmentation, linear interpolation for missing values, channel-independent processing for dimensional alignment, and standardized metadata. All sequences are stored in the ARROW format (Peltenburg et al., 2019) for efficient access. Further Details are provided in Appendix A.1.
4.2.2. Hierarchical Series Tree
It supports efficient indexing and retrieval in CRB by partitioning the database into domain-based groups and using $k$-means clustering. This hierarchical structure accelerates approximate nearest-neighbor search and supports dynamic updates, ensuring the index evolves incrementally over time. Details are provided in Appendix 7.
4.2.3. Hybrid and Hierarchical Time-series Retrieval